Suppose a random variable has the normal distribution such that X∼(82,4.8)X∼(82,4.8). What value of the random variable cuts the bottom 10%? Give your answer to 4 decimal places.
Answer #1

Given that,

X∼(\small \mu= 82, \small \sigma = 4.8)

Therefore,

\small \frac{X-\mu}{\sigma}\sim N(0,1)

\small Let \ Z=\frac{X-\mu}{\sigma}

i.e. \ P(Z\le z)=\Phi(z)

Let 'a' be the value which cuts the bottom 10%.

So, by problem,

P(X<a)=0.10 (10% = 0.10)

or, \ P(\frac{X-\mu}{\sigma}<\frac{a-82}{4.8})=0.10

or, \ \Phi(\frac{a-82}{4.8})=0.10=\Phi(-1.28)

or, \ \frac{a-82}{4.8}=-1.28

or, \ a=75.856

The value which cuts the bottom 10% is 75.856.                 (Answer)

z table :

Z 0.0 0.1 02 03 0.4 05 0.6 0.7 0.8 0.9 1.0 1.1 12 13 14 15 1.6 17 18 1.9 2.0 21 22 23 2.4 25 2.6 2.7 2.8 2.9 3.0 3.1 3.2 3.3